OpenStudy (anonymous):

A car travels at uniform acceleration over a distance of 440.0 m. Its final velocity after 20.0 s is 33.0 m/s. What is its initial velocity?

OpenStudy (anonymous):

s = 440m t = 20s v = 33m/s subtitude the value into formula \[s = \frac{ 1 }{ 2 } (u + v)t\] \[440 = \frac{ 1 }{ 2 } (u + 33)(20)\] u = 11m/s